Question 3 (3 points) A password is six to eight characters long. Each password must contain at least one numerical digit. How many passwords are there?
Answer #1

Answer :

Let P be the number of possible passwords and P6,P7,P8 be the number of passwords of length 6,7,8 respectively.
By the product rule
number of strings of 6 characters = 36^6
number of strings with no digits = 26^6

P6 = 36^6 - 26^6
= 2,176,782,336 - 308,915,776
= 1,867,866,560

P7 = 36^7 - 26^7
= 78,364,164,096 - 8,013,810,176
= 70,332,353,920

P8 = 36^8 - 26^8
= 2,821,109,907,456 - 208,827,064,576
= 70,332,353,920

P = P6+P7+P8
= 1,867,866,560+70,332,353,920+70,332,353,920
= 2,684,483,063,360